# Mergers all the way down: stellar collisions and kinematics of a dense hierarchically forming massive star cluster in a dwarf starburst

### v1

@@ 3.2.2 Velocity distribution
Next we inspect the velocity dispersion and velocity anisotropy profiles, as well as whether the cluster rotates. We compute the radial and tangential velocity distributions in spherical coordinates $v_{\mathrm{rad}}$ and $v_{\mathrm{tan}}=\sqrt{v_{\theta}^{2}+v_{\phi}^{2}}$ (where $v_{\theta},v_{\phi}$ are the angular velocity components) as well as using the line of sight (LOS) velocity in a ±$\pm 1$ pc slit perpendicular to the cluster rotation axis, $V_{\mathrm{LOS}}$. The slit is placed according to the angular momentum vector within the 3D half-mass radius. The velocities are binned radially in 0.1 dex bins and the innermost bin includes all stars within that radius. The corresponding velocity dispersions $\sigma_{\mathrm{rad}},\sigma_{\theta},\sigma_{\phi}$, and $\sigma_{\mathrm{tan}}=\sqrt{\sigma_{\theta}^{2}+\sigma_{\phi}^{2}}$ and total $\sigma=\sqrt{\sigma_{\mathrm{rad}}^{2}+\sigma_{\theta}^{2}+\sigma_{\phi}^{2}}$ are computed using the standard deviation of each velocity component in each bin. The cartesian coordinates are centred at the cluster centre of mass. Here we also mask out stars that have velocities greater than $\sim 70$ km s-1, in excess of the bulk velocity distribution in the cluster, to remove stars that are either escaping or in close binaries. Standard deviations for the profiles are computed by bootstrapping the stars 100 times, each time in a random orientation to include the effect of inclination on $V_{\mathrm{LOS}}$. We use the velocity anisotropy parameter $\beta$ to describe the ratio in tangential and radial dispersion (Binney & Tremaine, 2008) as
@@
Fig. 4 shows the evolution of the velocity distribution in the most massive cluster with time. The increase of the velocity dispersion toward the centre steepens with time as the central density of the cluster increases (see Fig. 3). The peak central value is $\sigma\sim 37$ km s-1. Once the cluster settles with one dominant core (as opposed to several merging sub-clumps), its peak rotation velocity shifts inwards. By the final snapshot the peak value reaches $V_{\mathrm{LOS}}\sim 7$ km s-1 and $v_{\mathrm{tan}}\sim 12$ km s-1 at a radius of 0.3similar-toabsent0.3$\sim$ 0.3∼ 0.30.40.40.40.4 pc. For comparison, GCs have similar or lower rotation velocities measured both in 1D and 3D (Kamann et al., 2018; Bianchini et al., 2018; Leitinger et al., 2025), consistent with their dynamically more evolved state. R136 also rotates at a rate of a few km s-1 (Hénault-Brunet et al., 2012), which is slower than our simulated cluster and consistent with its lower mass and previous results on star cluster dynamics in the griffin project (Lahén et al., 2020b).
@@
The rotational support $(V/\sigma)_{\mathrm{tot}}$, computed as the ratio between peak rotation velocity within $3R_{\mathrm{eff}}$ and the central velocity dispersion, is 0.3similar-toabsent0.3$\sim$ 0.3∼ 0.30.50.50.50.5 regardless of using the LOS velocity or the tangential rotational velocity. The value of $(V/\sigma)_{\mathrm{tot}}$ remained at a similar value during the final 1.3 Myr of the cluster assembly. With outward angular momentum transport, $(V/\sigma)_{\mathrm{tot}}$ is then expected to reduce later with time. The obtained $(V/\sigma)_{\mathrm{tot}}$ is in agreement with observations of GCs where typical values range between 0.1 and 0.6 in GCs that show statistically significant amounts of ordered rotation and that are dynamically young (relaxation time $>\mathrm{Gyr}$; e.g. Bianchini et al. 2018). Other definitions for $V/\sigma$, e.g. using the rotation amplitude over the total velocity dispersion observed in GCs in Leitinger et al. (2025), also show similar values and similar correlation with dynamical age as in Bianchini et al. (2018).
Figure 5: The radial velocity distribution of the stars in the most massive cluster divided in to the younger (blue) and older (red) population in the final snapshot. The panels show the stellar number density (a); 3D velocity dispersion (b); radial and tangential velocity dispersion (c); velocity anisotropy excluding ($\beta$, solid) and including ($\beta_{\mathrm{rot}}$, dashed) the bulk velocities (d); the projected LOS rotation velocity in a ±$\pm 1$ pc slit and the tangential rotation velocity (solid and dashed lines, respectively; e); and the corresponding $V/\sigma$ of the tangential and LOS velocity components (f). The vertical gray line in the panels indicates the projected $R_{\mathrm{eff}}$, and the blue and red vertical lines in panel a shows the values of $R_{\mathrm{eff}}$ for the young and old populations, respectively.
@@ 3.2.3 Kinematics of the age groups
Finally we investigate whether the hierarchical cluster assembly leaves an imprint in the shape and kinematics of the stars. We divide the stars into two age groups either younger or older than the mean stellar age of the cluster in the final snapshot. The groups have similar total masses within 7%. The old population formed mainly in sub-clumps that accreted hierarchically before the cluster transitioned into the centrally concentrated star formation mode. The younger population corresponds to the latter. The age groups do not signify generations, but simply the two $\sim 3$ Myr halves of the fairly Gaussian star formation history of this cluster that proceeded over a time span of $\sim 6$ Myr (see Fig. 7 in Lahén et al. 2024 for a similar star formation history). The kinematic properties are computed for these two populations in the same radial bins as was used for the whole stellar population.
Fig. 5 shows the number density and velocity profiles of the two age groups. The young component is slightly more centrally concentrated with $R_{\mathrm{eff}}\sim 0.56$ pc, compared to $R_{\mathrm{eff}}\sim 0.96$ pc of the old component and $R_{\mathrm{eff}}\sim 0.75$ pc of the entire cluster. The young component has a significantly higher rotation velocity compared to the old component. The peak rotation rate of the old population of a few km s-1 is consistent with the residual random angular momentum resulting from the hierarchical assembly process, based on initially non-rotating hierarchical N-body models of similar mass (Rantala et al. 2024; Rantala et al. in prep). The rotation velocity seen in Fig. 4 (panels c and d) is therefore mainly dominated by the younger population of stars that has not yet relaxed unlike the old population. The radially biased anisotropy profile seen in Fig. 4 (panel b) at the outer parts is dominated by the younger component as well.
If we do not reduce the bulk velocity and only use the root mean square velocities to compute $\beta_{\mathrm{rot}}$, the young population shows at intermediate radii ($0.1$ pc $<r<2$ pc) a strong tangential bias (down to $\sim-1$) compared to the fairly isotropic profile at $r\lesssim R_{\mathrm{eff}}$ shown in Fig. 5. This is simply a reflection of the rotation, that is encoded in bulk tangential motion. The radial bias in the outer radii of the young population is still retained even with bulk rotation included, as already shown in Fig. 4. The old population, on the other hand, remains isotropic when bulk rotation is not removed. The only exception is in the outermost parts where the $\beta_{\mathrm{rot}}$ of the old population shows tangential bias, i.e. bulk rotation.
@@
Our results confirm the picture drawn out in previous simulation of star formation in pre-existing massive star clusters (Bekki 2010; McKenzie & Bekki 2021; Lacchin et al. 2022) wherein the second formed population rotates significantly more than the first population. The previous simulations assumed a gradual initial rotation velocity for the first population, while here we self-consistently follow the formation and evolution of both two populations.
@@ 3.2.4 Shapes of the age groups
Fig. 6 shows the shape parameters for stars enclosed within increasing radii in the entire cluster and in the two age groups. The core of the cluster has a significant ellipticity (up to $\epsilon\sim 0.6$) and a mildly prolate ($T>0.66$) or triaxial ($0.660.33<T<$ 0.66) shape reflecting the star formation out of the disky internal gas structure. The cluster becomes increasingly triaxial or oblate ($T<0.33$) when stars at outer radii are included, but note that the overall shape becomes more spherical e.g. with $\epsilon\sim 0.2$ within $R_{\mathrm{eff}}$. The older population drives down the ellipticity in the cluster as it is fairly triaxial/round, while the young population has a more elliptic shape throughout the cluster.
Our new results verify the mildly elliptic shapes found for simulated young massive clusters in Lahén et al. (2020b), where collisional dynamics was not yet considered. Observed (evolved) GCs show similarly low ellipticities (Bianchini et al., 2018), hence we conclude that massive dense clusters may have such spherical shapes already at birth.
@@ 3.2.5 Comparison to NGC 104/47 Tuc
The most massive simulated young cluster has similar rotation characteristics with the Galactic GC NGC 104/47 Tuc that has been widely analysed in kinematic surveys (e.g. Lane et al. 2010; Bellini et al. 2017; Bianchini et al. 2018). NGC 104 is already $\gtrsim 11$ Gyr old (Gratton et al., 2003; Dotter et al., 2010), however it is still among the dynamically younger GCs (Baumgardt & Hilker, 2018). The differences in spatial and kinematic properties of our simulated young and old stellar populations (centrally more concentrated young stars; faster rotating young stars; radially biased outer young stars) are similar to what is found in some studies of NGC 104 that compare the chemically peculiar P2 and chemically normal P1 stars (P2 is more centrally concentrated; P2 rotates faster; P2 has a radial anisotropy in the outer parts; e.g. Milone et al. 2012; Dalessandro et al. 2024; Leitinger et al. 2025; but see also e.g. Martens et al. 2023 where no statistically significant difference between is found for P2 and P1). While the two age groups defined here are not directly analogous to the chemically identified multiple populations, the striking qualitative resemblance of the spatial and kinematic properties of the young and old population with the observed P2 and P1 in NGC 104 may signify a similar temporal sequence in their formation. In this scenario the P2 of NGC 104 would indeed have formed after P1 in a more centrally concentrated star formation phase.
It remains to be tested whether the simulated populations retain their differences over an extended period of time. Mastrobuono-Battisti & Perets (2013) and Hénault-Brunet et al. (2015) for instance showed with N-body simulations that a disk-like young population can retain the input lower velocity dispersion and outer radial anisotropy for a long time. The specific internal dynamical evolution depends on the initial stellar distribution and kinematics, as well as on the evolution of the tidal field. The present results would thus offer more self-consistent initial conditions for follow-up N-body simulations.
@@ 4 Conclusions and discussion
The most massive star cluster has a gravitationally bound mass of $\sim 2\times 10^{5}$ $\mathrm{M}_{\odot}$. This is similar to what was found in the previous collisionless simulation. The $\lesssim 1$ pc effective radius and surface density of $\sim 7\times 10^{4}$ $\mathrm{M}_{\odot}$ pc-2 in the most massive cluster approaches the extreme values of surface density and compactness ([$M_{\mathrm{cluster}}/10^{5}$ $\mathrm{M}_{\odot}$$]/[R_{\mathrm{eff}}/$ pc]>1]>1] > 1) of observed young massive star clusters found in recent high-redshift data (e.g. Adamo et al. 2024). The density profiles increase toward the cluster centre, without a flat core. The central density (central $>1000$ stars) is more than two orders of magnitude higher and the central surface density an order of magnitude higher compared to the effective values. This result is consistent with the resolved best-fit profiles of McLaughlin & van der Marel (2005) where it is predominantly the dynamically young massive clusters (relaxation time $>$Gyrs) that show central-to-effective surface density ratios ($\Sigma_{0}/\Sigma_{\mathrm{eff}}$) of more than 10. The implication is that the GC progenitors observed at high-redshifts probably had at least an order of magnitude higher central densities than the resolution limited mean values of $\Sigma_{\mathrm{eff}}\sim 10^{5}$$10^{6}$ $\mathrm{M}_{\odot}$ pc-2 reported thus far.
[…]
@@
The radial stellar surface density and surface brightness profiles of the most massive simulated cluster are fairly similar in shape with the most massive local young star cluster R136. The shape and $V/\sigma$ values agree with dynamically young GCs. The kinematic properties of the young and old populations are in qualitatively agreement with those of the chemically peculiar P2 and normal P1 stars, respectively, in the Galactic GC NGC 104. In some other dynamically young GCs, such as NGC 3201 and NGC 6101, the outer radii are also radially anisotropic, seemingly driven by the P2 stars (Leitinger et al., 2025). These clusters, however, show no clear differences in the rotation rates of the populations, and they have more centrally concentrated P1 populations as opposed to NGC 104 that shows a more centrally concentrated P2. Explaining the variety of population characteristics in GCs will thus require further investigation, including long-timescale N-body simulations.
@@
Our stellar collision product of $\sim 1000$ $\mathrm{M}_{\odot}$ would result in a black hole of at least $\sim 600$ $\mathrm{M}_{\odot}$ at our adopted metallicity of $0.01$ $\mathrm{Z}_{\odot}$. The most massive star and the resulting black hole are therefore $\sim 0.5\%$ of the total mass of the cluster. This is similar to the lower limits for the most massive stars found in N-body simulations where the collision products typically contain between a fraction of a per cent to $\sim 10\%$ of the cluster mass (Portegies Zwart et al., 2004; Arca Sedda et al., 2023; Rantala et al., 2024). Between a few per cent and $10\%$ is also the expected upper limit for extremely massive or supermassive stars forming through collisions or gas accretion based on analytic arguments (Gieles et al., 2018; Gieles et al., 2025). We are therefore starting to probe the very high mass end of the stellar mass function, often found theoretically to extend to $10^{3}$ $\mathrm{M}_{\odot}$ or $10^{4}$ $\mathrm{M}_{\odot}$ (e.g. Fujii et al. 2024; Chon & Omukai 2024), now in galactic scale populations of simulated star clusters.
@@
Our future aim is to explore the chemo-dynamical origin of multiple populations in GC progenitor clusters. While strong nitrogen enhancements are easily produced even with small amounts of hot hydrogen burning products, the currently implemented enrichment sources have not been able to produce large numbers of stars with as strong enhancements of, for instance, Al and Mg as observed (Lahén et al. 2024). This remains true even if we add the mass loss of up to $400$ $\mathrm{M}_{\odot}$ from collisions and VMS winds related to the growth and evolution of the most massive star in the presently most massive cluster. Additional mass growth, for instance, of the most massive star to the extreme or supermassive regime (Denissenkov & Hartwick, 2014; Gieles et al., 2025) will be needed to supply the star forming gas rapidly with significant amounts of light elements to explain the abundance anomalies in GCs. The growth could be enhanced through increased collision rates via primordial binaries (Fregeau et al., 2004; González Prieto et al., 2024; Rantala & Naab, 2025) or with sustained gas accretion and associated mass loss (Gieles et al., 2018). The stellar evolution methodology can also be expanded in the future with more accurate criteria and non-mass conserving prescriptions for mass transfer and collisions (e.g. adopting the Roche lobe overflow criterion from Eggleton 1983). The chemical enrichment due to TDEs (Kochanek, 2016) and binary evolution (de Mink et al., 2009; Nguyen & Sills, 2024) may be explored with a method that returns the lost mass back into the ISM.

### v2

@@ 3.2.2 Velocity distribution
Next we inspect the velocity dispersion and velocity anisotropy profiles, as well as whether the cluster rotates. We compute the radial and tangential velocity distributions in spherical coordinates $v_{\mathrm{rad}}$ and $v_{\mathrm{tan}}=\sqrt{v_{\theta}^{2}+v_{\phi}^{2}}$ (where $v_{\theta},v_{\phi}$ are the angular velocity components) as well as using the line of sight (LOS) velocity in a $\pm 1$ pc slit perpendicular to the cluster rotation axis, $V_{\mathrm{LOS}}$. The slit is placed according to the angular momentum vector within the 3D half-mass radius. The velocities are binned radially in 0.1 dex bins and the innermost bin includes all stars within that radius. The corresponding velocity dispersions $\sigma_{\mathrm{rad}},\sigma_{\theta},\sigma_{\phi}$, and $\sigma_{\mathrm{tan}}=\sqrt{\sigma_{\theta}^{2}+\sigma_{\phi}^{2}}$ and total $\sigma=\sqrt{\sigma_{\mathrm{rad}}^{2}+\sigma_{\theta}^{2}+\sigma_{\phi}^{2}}$ are computed using the standard deviation of each velocity component in each bin. The cartesian coordinates are centred at the cluster centre of mass. Here we also mask out stars that have velocities greater than $\sim 70$ km s-1, in excess of the bulk velocity distribution in the cluster, to remove stars that are either escaping or in close binaries. Standard deviations for the profiles are computed by bootstrapping the stars 100 times, each time in a random orientation to include the effect of inclination on $V_{\mathrm{LOS}}$. We use the velocity anisotropy parameter $\beta$ to describe the ratio in tangential and radial dispersion (Binney & Tremaine, 2008) as
@@
Fig. 4 shows the evolution of the velocity distribution in the most massive cluster with time. The increase of the velocity dispersion toward the centre steepens with time as the central density of the cluster increases (see Fig. 3). The peak central value is $\sigma\sim 37$ km s-1. Once the cluster settles with one dominant core (as opposed to several merging sub-clumps), its peak rotation velocity shifts inwards. By the final snapshot the peak value reaches $V_{\mathrm{LOS}}\sim 7$ km s-1 and $v_{\mathrm{tan}}\sim 12$ km s-1 at a radius of $\sim 0.3$$0.4$ pc. For comparison, GCs have similar or lower rotation velocities measured both in 1D and 3D (Kamann et al., 2018; Bianchini et al., 2018; Leitinger et al., 2025), consistent with their dynamically more evolved state. R136 also rotates at a rate of a few km s-1 (Hénault-Brunet et al., 2012), which is slower than our simulated cluster and consistent with its lower mass and previous results on star cluster dynamics in the griffin project (Lahén et al., 2020b).
@@
The rotational support $(V/\sigma)_{\mathrm{tot}}$, computed as the ratio between peak rotation velocity within $3R_{\mathrm{eff}}$ and the central velocity dispersion, is $\sim 0.3$$0.5$ regardless of using the LOS velocity or the tangential rotational velocity. The value of $(V/\sigma)_{\mathrm{tot}}$ remained at a similar value during the final 1.3 Myr of the cluster assembly. With outward angular momentum transport, $(V/\sigma)_{\mathrm{tot}}$ is then expected to reduce later with time. The obtained $(V/\sigma)_{\mathrm{tot}}$ is in agreement with observations of GCs where typical values range between 0.1 and 0.6 in GCs that show statistically significant amounts of ordered rotation and that are dynamically young (relaxation time $\mathrm{Gyr}$; e.g. Bianchini et al. 2018). Other definitions for $V/\sigma$, e.g. using the rotation amplitude over the total velocity dispersion observed in GCs in Leitinger et al. (2025), also show similar values and similar correlation with dynamical age as in Bianchini et al. (2018).
Figure 5: The radial velocity distribution of the stars in the most massive cluster divided in to the younger (blue) and older (red) population in the final snapshot. The panels show the stellar number density (a); 3D velocity dispersion (b); radial and tangential velocity dispersion (c); velocity anisotropy excluding ($\beta$, solid) and including ($\beta_{\mathrm{rot}}$, dashed) the bulk velocities (d); the projected LOS rotation velocity in a $\pm 1$ pc slit and the tangential rotation velocity (solid and dashed lines, respectively; e); and the corresponding $V/\sigma$ of the tangential and LOS velocity components (f). The vertical gray line in the panels indicates the projected $R_{\mathrm{eff}}$, and the blue and red vertical lines in panel a shows the values of $R_{\mathrm{eff}}$ for the young and old populations, respectively.
@@ 3.2.3 Kinematics of the age groups
Finally we investigate whether the hierarchical cluster assembly leaves an imprint in the shape and kinematics of the stars. We divide the stars into two age groups either younger or older than the mean stellar age of the cluster in the final snapshot. The groups of stars have similar total masses within 7%. The older population formed mainly in sub-clumps that accreted hierarchically before the cluster transitioned into the centrally concentrated star formation mode. The younger population corresponds to the latter. To draw a qualitative parallel to the MPs of GCs, we may suppose that the formation of the chemically peculiar P2 occurs after the first VMSs have had time to enrich their surrounding, and thus the youngest stars in the cluster should also be the most chemically enriched. The age groups do not signify generations, but simply the two $\sim 3$ Myr halves of the fairly Gaussian star formation history of this cluster that proceeded over a time span of $\sim 6$ Myr (see Fig. 7 in Lahén et al. 2024 for a similar star formation history).
We compute the kinematic properties for these two populations in the same radial bins used for the whole stellar population. Fig. 5 shows the number density and velocity profiles of the two age groups. The young component is slightly more centrally concentrated with $R_{\mathrm{eff}}\sim 0.56$ pc, compared to $R_{\mathrm{eff}}\sim 0.96$ pc of the old component and $R_{\mathrm{eff}}\sim 0.75$ pc of the entire cluster. The young component has a significantly higher rotation velocity compared to the old component. The peak rotation rate of the old population of a few km s-1 is consistent with the residual random angular momentum resulting from the hierarchical assembly process, based on initially non-rotating hierarchical N-body models of similar mass (Rantala et al. 2024; Rantala et al. 2025). The rotation velocity seen in Fig. 4 (panels c and d) is therefore mainly dominated by the younger population of stars that has not yet relaxed unlike the old population. The radially biased anisotropy profile seen in Fig. 4 (panel b) at the outer parts is dominated by the younger component as well.
If we do not reduce the bulk velocity and only use the root mean square velocities to compute $\beta_{\mathrm{rot}}$, the young population shows at intermediate radii ($0.1$ pc <r<2<r<2 pc) a strong tangential bias (down to $\sim-1$) compared to the fairly isotropic profile at $r\lesssim R_{\mathrm{eff}}$ shown in Fig. 5. This is simply a reflection of the rotation, that is encoded in bulk tangential motion. The radial bias in the outer radii of the young population is still retained even with bulk rotation included, as already shown in Fig. 4. The old population, on the other hand, remains isotropic when bulk rotation is not removed. The only exception is in the outermost parts where the $\beta_{\mathrm{rot}}$ of the old population shows tangential bias, i.e. bulk rotation.
@@
In the absence of strong chemical variations as expected of MPs, another way of differentiating the sub-populations in addition to age would be based on their birth location. As the most massive stars are located in the core of the cluster, then this would be expected as the location where the most chemically enriched stars form. Our previous simulations indicate that the chemical enrichment does indeed occur concentrated in the central regions of massive clusters (Lahén et al., 2024). We tested performing the above analysis using a division into two populations in which one (here the second population) contains stars born within the half-mass radius of the most massive progenitor sub-cluster. The other population consists of stars that originate from the outer regions of the progenitor or from accreted sub-clumps (here the first population). The results for Fig. 5 remain qualitatively similar. The peak rotation rate of the inner population is reduced by 1–2 km s-1 from that shown for the young population in panel $e$ of Fig. 5. The rotation rate of the outer population peaks with $\sim 5$ km s-1 at $\sim 1$ pc. Some outer radial anisotropy is introduced into the outer population compared to the mostly isotropic old stars in panel $d$ of Fig. 5. The most drastic change is, however, in the density profiles. The inner-formed population remains strongly centrally concentrated and dominates the density profile within $r_{\mathrm{eff}}$ even though the stars only constitute 1/3rd of the total cluster mass. To regain a less biased prediction of the density distributions, we chose to show only the populations divided by age in Fig. 5. Stellar age can be viewed as a proxy for the progress of chemical enrichment in the cluster, and only selecting by age we allow later-formed stars also in the accreted sub-clusters to be considered a part of the population that would be analogous to P2 in observed GCs.
With this, we conclude that our results confirm the picture drawn out in previous simulation of star formation in pre-existing massive star clusters (Bekki 2010; McKenzie & Bekki 2021; Lacchin et al. 2022) wherein the second formed population rotates significantly more than the first population. The previous simulations assumed a gradual initial rotation velocity for the first population, while here we self-consistently follow the formation and evolution of both two populations.
@@ 3.2.4 Shapes of the age groups
Fig. 6 shows the shape parameters for stars enclosed within increasing radii in the entire cluster and in the two age groups. The core of the cluster has a significant ellipticity (up to $\epsilon\sim 0.6$) and a mildly prolate ($T>0.66$) or triaxial ($0.33<T<0.66$) shape reflecting the star formation out of the disky internal gas structure. The cluster becomes increasingly triaxial or oblate ($T<0.33$) when stars at outer radii are included, but note that the overall shape becomes more spherical e.g. with $\epsilon\sim 0.2$ within $R_{\mathrm{eff}}$. The older population drives down the ellipticity in the cluster as it is fairly triaxial/round, while the young population has a more elliptic shape throughout the cluster.
Our new results verify the mildly elliptic shapes found for simulated young massive clusters in Lahén et al. (2020b), where collisional dynamics was not yet considered. Observed (evolved) GCs show similarly low ellipticities (Harris, 1996), hence we conclude that massive dense clusters may have such spherical shapes already at birth.
@@ 3.2.5 Comparison to NGC 104/47 Tuc
The most massive simulated young cluster has similar rotation characteristics with the Galactic GC NGC 104/47 Tuc that has been widely analysed in kinematic surveys (e.g. Lane et al. 2010; Bellini et al. 2017; Bianchini et al. 2018). NGC 104 is already $\gtrsim 11$ Gyr old (Gratton et al., 2003; Dotter et al., 2010), however it is still among the dynamically younger GCs (Baumgardt & Hilker, 2018). The differences in spatial and kinematic properties of our simulated young and old stellar populations (centrally more concentrated young stars; faster rotating young stars; radially biased outer young stars) are similar to what is found in some studies of NGC 104 that compare the chemically peculiar P2 and chemically normal P1 stars (P2 is more centrally concentrated; P2 rotates faster; P2 has a radial anisotropy in the outer parts; e.g. Milone et al. 2012; Dalessandro et al. 2024; Leitinger et al. 2025; but see also e.g. Martens et al. 2023 where no statistically significant difference between is found for P2 and P1). While the two age groups defined here are not directly analogous to the chemically identified multiple populations, the striking qualitative resemblance of the spatial and kinematic properties of the young and old population with the observed P2 and P1 in NGC 104 may signify a similar temporal sequence in their formation. In this scenario the P2 of NGC 104 would indeed have formed after P1 in a more centrally concentrated star formation phase. Previous cloud-scale simulations have also suggested that massive stars would form only late in the cluster assembly process (Vázquez-Semadeni et al., 2019) supporting the late formation of the chemically peculiar population.
In some of the other dynamically young GCs, such as NGC 3201 and NGC 6101, the outer radii are also radially anisotropic, seemingly driven by the P2 stars (Leitinger et al., 2025). These clusters, however, show no clear differences in the rotation rates of the populations, and they have more centrally concentrated P1 populations as opposed to the more centrally concentrated P2 of NGC 104.
It remains to be tested whether the simulated populations retain their differences over an extended period of time. Mastrobuono-Battisti & Perets (2013) and Hénault-Brunet et al. (2015) for instance showed with N-body simulations that a disk-like young population can retain the input lower velocity dispersion and outer radial anisotropy for a long time. The specific internal dynamical evolution depends on the initial stellar distribution and kinematics, as well as on the evolution of the tidal field (e.g. Berczik et al. 2025). The present results would thus offer more self-consistent initial conditions for follow-up N-body simulations.
@@ 4 Conclusions and discussion
The most massive star cluster has a gravitationally bound mass of $\sim 2\times 10^{5}$ $\mathrm{M}_{\odot}$. This is similar to what was found in the previous collisionless simulation. The $\lesssim 1$ pc effective radius and surface density of $\sim 7\times 10^{4}$ $\mathrm{M}_{\odot}$ pc-2 in the most massive cluster approaches the extreme values of surface density and compactness ($_{\mathrm{cluster}}/10^{5} M\mathrm{M}_{\odot}$$R_{\mathrm{eff}}/$ pc]>1]>1) of observed young massive star clusters found in recent high-redshift data (e.g. Adamo et al. 2024). The density profiles increase toward the cluster centre, without a flat core. The central density (central >1000>1000 stars) is more than two orders of magnitude higher and the central surface density an order of magnitude higher compared to the effective values. This result is consistent with the resolved best-fit profiles of McLaughlin & van der Marel (2005) where it is predominantly the dynamically young massive clusters (relaxation time >>Gyrs) that show central-to-effective surface density ratios ($\Sigma_{0}/\Sigma_{\mathrm{eff}}$) of more than 10. The implication is that the GC progenitors observed at high-redshifts probably had at least an order of magnitude higher central densities than the resolution limited mean values of $\Sigma_{\mathrm{eff}}\sim 10^{5}$$10^{6}$ $\mathrm{M}_{\odot}$ pc-2 reported thus far.
The most massive star cluster rotates at a rate of $V_{\mathrm{LOS}}\sim$7 km s-1 or $v_{\mathrm{tan}}\sim$12 km s-1, with the radial rotation profile peaking inside the half-mass radius. To decipher which stars dominate the rotation, we further divided the stars of the cluster into a younger and an older half across the $\sim 6$ Myr of star formation in the cluster. We find that the younger population is more centrally concentrated, it has a significantly faster peak rotation velocity (>12>12 km s-1 compared to 2–4 km s-1 in the old population), and its velocity distribution is radially anisotropic in the outer parts. The first formed, older population is more isotropic and it has a larger velocity dispersion. The cluster has a significant elliptic and mildly prolate shape in the inner parts while the overall shape of the cluster is mildly elliptic ($\epsilon\sim 0.1$$0.3$). The older population is more triaxial/round in shape throughout the cluster while the young population has a milddy prolate shape in the inner parts and an oblate shape in the outer parts. The first formed population has thus already relaxed throughout the hierarchical assembly process, while the younger population exhibits its initial shape and kinematics, including some radial motions in the outer parts from sub-clump infall.
@@
The radial stellar surface density and surface brightness profiles of the most massive simulated cluster are fairly similar in shape with the most massive local young star cluster R136. The shape and $V/\sigma$ values agree with dynamically young GCs. The kinematic properties of the young and old populations are in qualitatively agreement with those of the chemically peculiar P2 and normal P1 stars, respectively, in the Galactic GC NGC 104. In other dynamically young clusters, such as NGC 3201 and NGC 6101, the differences in the spatio-kinematics of the stellar populations can be less pronounced or even reversed compared to the features found here (e.g. a centrally more concentrated P1). Explaining the variety of population characteristics in GCs will thus require further investigation, including long-timescale N-body simulations.
@@
Our stellar collision product of $\sim 1000$ $\mathrm{M}_{\odot}$ would result in a black hole of at least $\sim 600$ $\mathrm{M}_{\odot}$ at our adopted metallicity of $0.01$ $\mathrm{Z}_{\odot}$. The most massive star and the resulting black hole are therefore $\sim 0.5\%$ of the total mass of the cluster. This is similar to the lower limits for the most massive stars found in N-body simulations where the collision products typically contain between a fraction of a per cent to $\sim 10\%$ of the cluster mass (Portegies Zwart et al., 2004; Arca Sedda et al., 2023; Rantala et al., 2024; Rantala et al., 2025). Between a few per cent and $10\%$ is also the expected upper limit for extremely massive or supermassive stars forming through collisions or gas accretion based on analytic arguments (Gieles et al., 2018; Gieles et al., 2025). We are therefore starting to probe the very high mass end of the stellar mass function, often found theoretically to extend to $10^{3}$ $\mathrm{M}_{\odot}$ or $10^{4}$ $\mathrm{M}_{\odot}$ (e.g. Fujii et al. 2024; Chon & Omukai 2025), now in galactic scale populations of simulated star clusters.
@@
Our future aim is to explore the chemo-dynamical origin of multiple populations in GC progenitor clusters. While strong nitrogen enhancements are easily produced even with small amounts of hot hydrogen burning products, the currently implemented enrichment sources have not been able to produce large numbers of stars with as strong enhancements of, for instance, Al and Mg as observed (Lahén et al. 2024). This remains true even if we add the mass loss of up to $400$ $\mathrm{M}_{\odot}$ from collisions and VMS winds related to the growth and evolution of the most massive star in the presently most massive cluster. Additional mass loss, for instance, through disruption of stars (Elmegreen, 2017) or the growth of the most massive star to the extreme or supermassive regime (Denissenkov & Hartwick, 2014; Gieles et al., 2025) will be needed to supply the star forming gas rapidly with significant amounts of light elements to explain the abundance anomalies in GCs. The growth could be enhanced through increased collision rates via primordial binaries (Fregeau et al., 2004; González Prieto et al., 2024; Rantala & Naab, 2025) or with sustained gas accretion and associated mass loss (Gieles et al., 2018; Ramírez-Galeano et al., 2025). The stellar evolution methodology can also be expanded in the future with more accurate criteria and non-mass conserving prescriptions for mass transfer and collisions (e.g. adopting the Roche lobe overflow criterion from Eggleton 1983). The chemical enrichment due to TDEs (Kochanek, 2016) and binary evolution (de Mink et al., 2009; Nguyen & Sills, 2024) may be explored with a method that returns the lost mass back into the ISM.